# ROTI-GCV: Generalized Cross-Validation for right-ROTationally Invariant Data

### v1

@@ Abstract
Two key tasks in high-dimensional regularized regression are tuning the regularization strength for good predictions and estimating the out-of-sample risk. It is known that the standard approach — $k$-fold cross-validation — is inconsistent in modern high-dimensional settings. While leave-one-out and generalized cross-validation remain consistent in some high-dimensional cases, they become inconsistent when samples are dependent or contain heavy-tailed covariates. To model structured sample dependence and heavy tails, we use right-rotationally invariant covariate distributions - a crucial concept from compressed sensing. In the common modern proportional asymptotics regime where the number of features and samples grow comparably, we introduce a new framework, ROTI-GCV, for reliably performing cross-validation. Along the way, we propose new estimators for the signal-to-noise ratio and noise variance under these challenging conditions. We conduct extensive experiments that demonstrate the power of our approach and its superiority over existing methods.
@@ 1 Introduction
Tuning the regularization strength in penalized regression estimators is a fundamental technique for addressing overfitting in machine learning and data science, particularly important for modern high-dimensional datasets. In this work, we propose a new cross-validation technique, ROTI-GCV, for determining one of the most fundamental hyperparameters, the ridge regularization parameter $\lambda$, under a different model on the design $\mathbf{X}$. The classical approach for tuning $\lambda$ is $k$-fold cross-validation [20], where the training data is divided into $k$ equally sized chunks (folds). A model is then repeatedly trained on all but one fold and validated on the remaining fold. The value of $\lambda$ achieving the best average validation performance is then selected. However, in high-dimensional settings where the dataset dimension $p$ scales proportionally with dataset size $n$, using a finite $k$ results in an inconsistent estimate of out-of-sample error [32, 43]. On the other hand, leave-one-out cross-validation (LOOCV, corresponding to $k=n$) and the lesser-known generalized cross-validation (GCV) remain consistent in some high-dimensional settings [19, 3, 31].
The theoretical consistency of LOOCV and GCV in high dimensions has thus far been based on the assumption that the samples are independent and identically distributed (i.i.d.) draws from some fixed distribution of sufficient regularity, such as having light enough tails [20, 3, 31]. These assumptions, however, do not always hold in practice. We show in this study that in settings where there is dependence across observations or that the data is heavy-tailed, LOOCV and GCV become biased. To address the limitations of i.i.d. assumptions, we propose modeling designs using an alternative random matrix ensemble. Instead of assuming the rows of our design to be i.i.d. draws from a fixed distribution, we instead require that the singular value decomposition of the design $\mathbf{X}$ is right-rotationally invariant (see Definition 1).
Right-rotationally invariant designs form a rich class, allowing for tractable analysis while also capturing structured forms of dependence among datapoints. These designs have recently received much attention as alternatives to fully i.i.d. designs for theoretical analysis in a multitude of works [38, 29, 33, 39, 7, 11, 14, 40, 41, 42, 15, 28, 26, 46, 27, 13, 12]. Recent works [8, 44] also show that, under certain conditions, the behavior of high dimensional matrices is largely dictated by their spectrum, and in fact will “act” similarly to a right-rotationally invariant matrices with the same spectrum.
In this work, we introduce a new framework inspired by generalized cross-validation, which we call ROTI-GCV, that is provably consistent for the out-of-sample risk when the data is right-rotationally invariant. We then provide simulations demonstrating that the new formulation outperforms traditional GCV and LOOCV in scenarios where observations are correlated and exhibit heavy tails.
@@ 2.1 Right-rotationally invariant designs
A random design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ is right-rotationally invariant if for any $\mathbf{O}\in\mathbb{O}(p)$, one has $\mathbf{XO}\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}$, where $\mathbb{O}(p)$ denotes the group of $p\times p$ orthogonal matrices. Equivalently, let $\mathbf{X}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}$ be the singular value decomposition of $\mathbf{X}$. Then $\mathbf{X}$ is right-rotationally invariant if and only if $\mathbf{O}$ is independent of $(\mathbf{Q},\mathbf{D})$ and $\mathbf{O}$ drawn from the Haar measure on $\mathbb{O}(p)$ that is, it is uniformly distributed on $\mathbb{O}(p)$.
The setting where each entry of $\mathbf{X}$ is an i.i.d. standard Gaussian is the simplest example of a right-rotationally invariant design. Other examples of right-rotationally invariant designs include the following:
  • Equicorrelated data: $\mathbf{X}\in\mathbb{R}^{n\times p}$ has independent columns, but each column follows a multivariate Gaussian distribution with covariance matrix $\boldsymbol{\Sigma}$, where $\Sigma_{ij}=\rho$ if $i\neq j$, and $\Sigma_{ii}=1$. This can model situations where the training dataset a populations sharing a latent factor (e.g. location), causing the samples to have correlated covariates.
  • Autocorrelated data: The rows of $\mathbf{X}$ can be drawn from a time series, where $\mathbf{x}_{i}=\rho\mathbf{x}_{i-1}+\sqrt{1-\rho^{2}}\boldsymbol{\epsilon}_{i}$, with $\boldsymbol{\epsilon}_{i}$ being i.i.d. draws from $\mathsf{N}(0,\mathbf{I}_{n})$.
@@
  • spiked matrices: $\mathbf{X}=\lambda\mathbf{V}\mathbf{W}^{\top}+\mathbf{G}$, where $\mathbf{V}\in\mathbb{R}^{n\times r}$ and $\mathbf{W}\in\mathbb{R}^{p\times r}$ are the first $r$ columns of two Haar matrices.
See [27, Figure 1] for more such examples.
In essence, right-rotationally invariant ensembles allow us enormous flexibility—they can simultaneously capture dependent rows and heavy-tails by allowing a rather general class of singular value distributions. The spectrum of right-rotationally invariant designs can essentially be arbitrary, and allowing for significant deviations from prior i.i.d. designs used to study the behavior of LOOCV and GCV in high dimensions. Furthermore, their universality class is extremely broad, as established in [8], capturing even challenging deterministic designs arising in compressed sensing. At the same time, the fact that $\mathbf{O}$ is Haar does add some restrictions - in particular, the designs we consider are now necessarily isotropic.
Lemma 1.
In prior analyses, the spectrum of $\boldsymbol{\Sigma}$ defines the problem geometry. We sacrifice this degree of freedom for the ability to set the spectrum $\mathbf{D}$ to capture distributions beyond those of i.i.d. rows.
To simplify some statements, we introduce some notation. Let $m(\mu,z)$ denote the Stieltjes transform of a measure $\mu$ on $\mathbb{R}$. In our analysis, we will mostly care when $\mu$ is the empirical distribution of the entries of $\mathbf{D}^{\top}\mathbf{D}$ (which are the same as those of $\mathbf{X}^{\top}\mathbf{X}$). Take $\mu_{\mathbf{D}^{\top}\mathbf{D}}$ to denote the empirical measure of the diagonal entries of $\mathbf{D}^{\top}\mathbf{D}$, and let $m_{\mathbf{D}}(z)=m(\mu_{\mathbf{D}^{\top}\mathbf{D}},z)$. We also consider the companion Stieltjes transform, which is the Stieltjes transform of the empirical distribution of $\mathbf{D}\mathbf{D}^{\top}$. We will denote this as $v_{\mathbf{D}}(z)=m(\mu_{\mathbf{D}\mathbf{D}^{\top}},z)$.
@@ 2.2 Problem Setup
We study high-dimensional linear regression with right-rotationally invariant design. Concretely, one observes $(\mathbf{X},\mathbf{y})$, such that
@@
with $\mathbf{y}\in\mathbb{R}^{n}$, $\mathbf{X}\in\mathbb{R}^{n\times p}$, and $\boldsymbol{\epsilon},\boldsymbol{\beta}\in\mathbb{R}^{p}$. Hence $n$ is the sample size, $p$ is both the data dimension and the parameter count, $\boldsymbol{\beta}$ is the signal, and $\boldsymbol{\epsilon}$ is a component-wise independent noise vector, where each entry has mean 00 and variance $\sigma^{2}$. We model $\mathbf{X}$ as being right-rotationally invariant, $\boldsymbol{\beta}$ as being fixed or random independent of $\mathbf{X}$, and $\boldsymbol{\epsilon}$ independent of everything else.
The problem we consider will be the expected out-of-sample risk on an independent draw from a new, potentially different right-rotationally invariant design, $\widetilde{\mathbf{X}}=\widetilde{\mathbf{Q}}^{\top}\widetilde{\mathbf{D}}\widetilde{\mathbf{O}}$, with $\widetilde{\mathbf{X}}\in\mathbb{R}^{n^{\prime}\times p}$. This corresponds, for example, to the setting of learning a model on one interdependent population (such as medical data from some population), and using the model on a separate interdependent population. The explicit risk we examine, following [31], is
@@
This setting is comparable to the isotropic setting in [19].
We study the risk of the ridge estimator
$\hat{\boldsymbol{\beta}}_{\lambda}=\operatorname*{arg\;min}_{\mathbf{b}\in\mathbb{R}^{p}}\left\{\|\mathbf{y}-\mathbf{X}\mathbf{b}\|_{2}^{2}+\lambda\|\mathbf{b}\|^{2}_{2}\right\}$ (3)
which has closed form $\hat{\boldsymbol{\beta}}=(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{y}$.
2.2.1 Scaling and Intuition
Lemma 1 shows that $\mathbb{E}[\mathbf{X}^{\top}\mathbf{X}]=\mathcal{O}(1)\cdot\mathbf{I}$, while in the i.i.d. setting with identity covariance, one would expect $\mathbb{E}[\mathbf{X}^{\top}\mathbf{X}/n]=\mathbf{I}$. Thus, $\mathbf{X}$ in our setting is akin to the normalized matrix $\mathbf{X}/\sqrt{n}$ in the i.i.d. setting. Furthermore, $\mathbf{D}^{\top}\mathbf{D}$ contains the eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$, and this scaling ensures that $\|\mathbf{D}^{\top}\mathbf{D}\|_{\mathsf{op}}=\mathcal{O}(1)$ in the setting where entries are i.i.d. draws from some fixed distribution. The consequence of this scaling is that in order to maintain the signal to noise ratio, we must have $\boldsymbol{\beta}$ having norm on the order of $\sqrt{n}$.

### v2

@@ Abstract
Two key tasks in high-dimensional regularized regression are tuning the regularization strength for accurate predictions and estimating the out-of-sample risk. It is known that the standard approach — $k$-fold cross-validation — is inconsistent in modern high-dimensional settings. While leave-one-out and generalized cross-validation remain consistent in some high-dimensional cases, they become inconsistent when samples are dependent or contain heavy-tailed covariates. As a first step towards modeling structured sample dependence and heavy tails, we use right-rotationally invariant covariate distributions — a crucial concept from compressed sensing. In the proportional asymptotics regime where the number of features and samples grow comparably, which is known to better reflect the empirical behavior in moderately sized datasets, we introduce a new framework, ROTI-GCV, for reliably performing cross-validation under these challenging conditions. Along the way, we propose new estimators for the signal-to-noise ratio and noise variance. We conduct experiments that demonstrate the accuracy of our approach in a variety of synthetic and semi-synthetic settings.
@@ 1 Introduction
Regularized estimators are a fundamental technique for addressing overfitting in machine learning and data science, particularly in modern high-dimensional datasets; thus, understanding their theoretical behavior has become all the more important. To this end, the prototypical problem of ridge regression has gained renewed attention, with multiple works characterizing its out-of-sample risk in high dimensions, c.f., [7, 18]. Additionally, consistent estimates for this out-of-sample risk, crucial for optimizing the regularization parameter $\lambda$, have been established in the form of leave-one-out cross-validation (LOOCV) and generalized cross-validation (GCV) [45, 29, 3].
Crucially, all aforementioned analyses assume that samples are independent and identically distributed (i.i.d.) draws from some distribution with sufficient regularity, such as light tails. These assumptions often fail in practice, such as in financial returns, neuroscience, and climate data [14, 48, 37]. In such cases, cross-validation techniques can become biased (Theorem 1, Remark 3), necessitating alternatives. Only recently, [4] considered independent but non-identically distributed samples and [2] explored Gaussian designs with both sample and feature dependence. Understanding cross-validation in high dimensions under more general sample dependencies and for designs with heavier tails than sub-Gaussians remains a significant challenge. This paper takes the first step in addressing this issue.
We assess the impact of sample dependence and heavy tails in ridge regression and cross-validation using an alternative random matrix ensemble for the covariate distribution. Instead of assuming a Gaussian matrix or i.i.d. rows from a fixed distribution,
we require the singular value decomposition of the design $\mathbf{X}$ to be right-rotationally invariant (see Definition 1). We characterize the ridge regression risk for these designs and introduce ROTI-GCV, a new GCV-inspired framework for consistently estimating the out-of-sample risk. Simulations show that ROTI-GCV outperforms traditional GCV and LOOCV in situations with correlated and heavy-tailed observations.
Right-rotationally invariant designs facilitate tractable analysis while capturing structured forms of sample dependencies and heavy-tailed covariates.
These designs have gained significant attention as alternatives to i.i.d. designs for theoretical analysis in numerous studies
[38, 26, 31, 9, 12, 39, 25, 22, 10, 46]. Recent work [8, 41] showed that, under mild conditions, properties of convex estimators across various designs are well-approximated by those under a right-rotationally invariant design with the same spectrum, indicating a broad universality class for these distributions.
Formally, we operate in the proportional asymptotics regime, where the dimension $p$ scales proportionally with the sample size $n$. This regime, rooted in probability theory and statistical physics, has received significant attention due to its ability to accurately capture empirical phenomena [18, 35, 36, 24]. Furthermore, results derived under this regime require minimal assumptions on the signal structure, resulting in theory and methods with broad practical utility [34]. In regards to cross-validation for i.i.d. samples, [18, 30, 40] established that classical $k$-fold cross-validation is inconsistent estimates for the out-of-sample error in this high-dimensional regime, whereas LOOCV and GCV remain consistent, prompting our study of GCV. However, our work focuses on non-i.i.d. samples and non-Gaussian tailed designs, areas not sufficiently addressed in prior work.
We provide proofs for ridge regression, but emphasize that our core idea can extend to other penalties. Our proof hinges on characterizing the limit of GCV (Theorem 1). This reveals that the original GCV is inconsistent for the out-of-sample risk in our setting. We next observe that though GCV is inconsistent, its asymptotic limit can serve as an estimating equation for the unknown parameters parameterizing the risk. This estimating equation approach directly inspires a new cross-validation method that works for our challenging right-rotationally invariant designs. A similar strategy may be invoked for any other penalty where an analogue of Theorem 1 is available. Therefore, we anticipate our approach to be broadly generalizable beyond ridge regression.
@@ 2.1 Right-rotationally invariant designs
A random design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ is right-rotationally invariant if for any $\mathbf{O}\in\mathbb{O}(p)$, one has $\mathbf{XO}\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}$, where $\mathbb{O}(p)$ denotes $p\times p$ orthogonal matrices. Equivalently, $\mathbf{X}$ is right-rotationally invariant if and only if the singular value decomposition $\mathbf{X}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}$ satisfies that $\mathbf{O}$ is independent of $(\mathbf{Q},\mathbf{D})$ and drawn from the Haar measure on $\mathbb{O}(p)$, i.e. “uniformly" distributed on $\mathbb{O}(p)$.
Some examples of right-rotationally invariant designs are as follows (note that an i.i.d. entries standard Gaussian design is a member of this class, but the class includes many non-i.i.d. and non-Gaussian designs as well; see Appendix B.3 for proofs of invariance):
  1. (i)
    Autocorrelated data: The rows of $\mathbf{X}$ can be drawn from an autocorrelated series, where $\mathbf{x}_{i}=\rho\mathbf{x}_{i-1}+\sqrt{1-\rho^{2}}\mathbf{z}_{i}$, with $\mathbf{z}_{i}$ being i.i.d. draws from $\mathsf{N}(0,\mathbf{I}_{n})$.
  2. (ii)
@@
  4. (iv)
    Matrix-Normals: $\mathbf{X}\sim\mathsf{MN}(\mathbf{0},\boldsymbol{\Sigma}^{\mathrm{row}},\boldsymbol{\Sigma}^{\mathrm{col}})$, where $\boldsymbol{\Sigma}^{\mathrm{row}}$ can be arbitrary while $\boldsymbol{\Sigma}^{\mathrm{col}}\sim\mathsf{InvWishart}(\mathbf{I}_{p},(1+\delta)p)$, for any $\delta>0$. See Appendix B.5 for details on distribution notation.
  5. (v)
    Equicorrelated data: $\mathbf{X}\in\mathbb{R}^{n\times p}$ has independent columns, but each column follows a multivariate Gaussian distribution with covariance matrix $\boldsymbol{\Sigma}$, where $\Sigma_{ij}=\rho$ if $i\neq j$, and $\Sigma_{ii}=1$.
  6. (vi)
    Spiked matrices: $\mathbf{X}=\lambda\mathbf{V}\mathbf{W}^{\top}+\mathbf{G}$, where $\mathbf{V}\in\mathbb{R}^{n\times r}$ and $\mathbf{W}\in\mathbb{R}^{p\times r}$ are the first $r$ columns of two Haar matrices, and $G_{ij}\sim\mathsf{N}(0,1)$.
See [23, Figure 1] for more such examples. Setting (ii) has heavier tails than supported by existing GCV theory; settings (i, iii, v, vi) have sample dependence; setting (iv) has both.
Right-rotationally invariant ensembles allow us enormous flexibility—they can simultaneously capture dependent rows and heavy-tails by allowing a rather general class of singular value distributions.
The spectrum of right-rotationally invariant designs can essentially be arbitrary, and allowing for significant generalization from prior i.i.d. designs used to study the behavior of LOOCV and GCV in high dimensions. Furthermore, the universality class of these designs is extremely broad, as established in [8, 41].
To simplify some statements, we introduce some notation. Let $m_{\mathbf{D}}(z)=\frac{1}{p}\sum_{i=1}^{p}\frac{1}{D_{ii}^{2}-z}$ and $v_{\mathbf{D}}(z)=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{D_{ii}^{2}-z}$, with $D_{ii}=0$ if $i>\min(n,p)$.
Also of use will be the derivatives $m^{\prime}_{\mathbf{D}}$ and $v^{\prime}_{\mathbf{D}}$. Note $m_{\mathbf{D}}$ is the Stieltjes transform of the empirical measure of $(\mathbf{D}^{\top}\mathbf{D}_{ii})_{i=1}^{p}$; $v_{\mathbf{D}}$ is the companion Stieltjes transform.
@@ 2.2 Problem Setup
We study high-dimensional ridge regression and GCV with right-rotationally invariant designs. Concretely, we observe $(\mathbf{X},\mathbf{y})$, such that
@@
with $\mathbf{y}\in\mathbb{R}^{n}$, $\mathbf{X}\in\mathbb{R}^{n\times p}$, and $\boldsymbol{\epsilon},\boldsymbol{\beta}\in\mathbb{R}^{p}$. Hence $n$ is the sample size, $p$ is both the data dimension and the parameter count, $\boldsymbol{\beta}$ is the signal, and $\boldsymbol{\epsilon}$ is a component-wise independent noise vector, where each entry has mean 00 and variance $\sigma^{2}$. We study the expected out-of-sample risk on an independent draw from a new, potentially different right-rotationally invariant design, $\widetilde{\mathbf{X}}=\widetilde{\mathbf{Q}}^{\top}\widetilde{\mathbf{D}}\widetilde{\mathbf{O}}$, with $\widetilde{\mathbf{X}}\in\mathbb{R}^{n^{\prime}\times p}$. This corresponds to the setting of learning a model on one interdependent population, and using the model on a separate interdependent population. We analyze the conditional risk
@@
Such conditional risks have been examined in prior works [18, 29].
We focus on the risk of the ridge estimator
$\hat{\boldsymbol{\beta}}_{\lambda}=\operatorname*{arg\;min}_{\mathbf{b}\in\mathbb{R}^{p}}\left\{\|\mathbf{y}-\mathbf{X}\mathbf{b}\|_{2}^{2}+\lambda\|\mathbf{b}\|^{2}_{2}\right\},$ (3)
which has closed form $\hat{\boldsymbol{\beta}}=(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{y}$. As such, sometimes we will write $R_{\mathbf{X},\mathbf{y}}(\lambda):=R_{\mathbf{X},\mathbf{y}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ for convenience. Before stating our results, we note the running assumptions used in what follows.
  1. A1
    We consider a sequence of right-rotationally invariant designs $\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{n},\dots$ where each $\mathbf{X}_{i}\sim\mathbf{Q}_{i}^{\top}\mathbf{D}_{i}\mathbf{O}_{i}$, and a sequence of signal vectors $\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2},\dots,\boldsymbol{\beta}_{n},\dots$ such that for all $n$, $\boldsymbol{\beta}_{n}\in\mathbb{R}^{p}$ and $\|\boldsymbol{\beta}_{n}\|=r\sqrt{n}$; $\mathbf{y}_{i}$ is generated by Eq. 1. In the sequel, we drop the dependence on $n$ and write $\mathbf{X},\mathbf{y}$ and $\boldsymbol{\beta}$ with the indexing implicit. We refer to $(\mathbf{X},\mathbf{y})$ as the training data.
  2. A2
    lim $\limsup\lambda_{\max}(\mathbf{D}_{n})<C$ a.s. for some constant $C$.
  3. A3
    $\mathbf{X}_{n}\in\mathbb{R}^{n\times p(n)}$, $n\to\infty$, $p(n)/n\rightarrow\gamma\in(0,\infty)$.
  4. A4
    $\boldsymbol{\epsilon}$ is independent of $\mathbf{X}$; $\epsilon_{i}$’s are independent; $\mathbb{E}[\epsilon_{i}]=0$, $\mathbb{E}[\epsilon_{i}^{2}]=\sigma^{2}$, and $\mathbb{E}[\epsilon_{i}^{4+\eta}]<\infty$ for some $\eta>0$.
  5. A5
    To fix scaling and simplify our results, we assume that $\frac{n\mathbb{E}[\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})]}{n^{\prime}p}=1$; this ensures the risk is defined and maintains the scale of the training set relative to the test set (See Remark 8 of Appendix).
  6. A6
    $\boldsymbol{\beta}_{n}$ is fixed or random independent of $(\mathbf{X}_{n},\boldsymbol{\epsilon})$.
  7. A7
    $\widetilde{\mathbf{X}}$ is an independent draw from a potentially different right-rotationally invariant ensemble, meaning $\widetilde{\mathbf{X}}\sim\widetilde{\mathbf{Q}}^{\top}\widetilde{\mathbf{D}}\widetilde{\mathbf{O}}$, with $(\widetilde{\mathbf{Q}},\widetilde{\mathbf{D}},\widetilde{\mathbf{O}})$ independent of $\mathbf{X}$. $\widetilde{\mathbf{Q}}$ and $\widetilde{\mathbf{D}}$ need not be distributed the same as $\mathbf{Q},\mathbf{D}$; $\widetilde{\mathbf{O}}$ must be Haar.
Of crucial importance will be the two parameters $r^{2}$ and $\sigma^{2}$, which dictate the behavior of the problem.
Remark 1.
To clarify why the risk in Eq. 2 is relevant, if one samples $(\widetilde{\mathbf{X}},\widetilde{\mathbf{y}})$ according to Eq. (1), then $R_{\mathbf{X},\mathbf{y}}(\lambda)=\mathbb{E}[\|\widetilde{\mathbf{y}}-\widetilde{\mathbf{X}}\hat{\boldsymbol{\beta}}_{\lambda}\|^{2}\mid\mathbf{X},\mathbf{y}]-\sigma^{2}$, meaning the risk defined corresponds, up to a factor of $\sigma^{2}$, to the mean-squared error on the test set.
Remark 2.
One should think of $\mathbf{X}$ in our setting as akin to the normalized matrix $\mathbf{X}/\sqrt{n}$ in the i.i.d. setting. This further ensures that $\|\mathbf{X}^{\top}\mathbf{X}\|_{\mathsf{op}}=\|\mathbf{D}^{\top}\mathbf{D}\|_{\mathsf{op}}=\mathcal{O}(1)$. To maintain the signal to noise ratio under this scaling, $\|\boldsymbol{\beta}\|=\mathcal{O}(\sqrt{n})$ (Assumption A1); the test set is appropriately scaled through Assumption A5.